Find an equation of a parabola with focus(0,-7) and directrix x = 7

Thanks so much in advance:)

Answer by TimothyLamb(4379)

---
given:
focus(0,-7)
directrix x = 7
---
directrix is a vertical line, so parabola is horizontal.
---
find vertex, half way between focus and directrix:
---
h = (7 - 0)/2 + 0
h = 7/2 = 3.5
vertex = (h,k) = (3.5, -7)
---
directrix is to the right of the focus so parabola opens to the left and p is negative
distance between focus and vertex: p = -3.5
---
conic form of horizontal parabola:
4p(x – h) = (y – k)^2
4(-3.5)(x – 3.5) = (y + 7)^2
---
answer:
-14(x – 3.5) = (y + 7)^2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php